OpenStudy (wholock):

I really need help with this. I don't need help solving the entire problem, just in figuring out in what order. A jet plane starts from rest on a runway. It accelerates uniformly at a rate of 20m/s^2. The equation for computing the distance traveled is d = (1/2)at^2. a. Find the distance traveled at the end of each second for five seconds. I know this is a function, but I need to know if I should apply the exponent first or multiply (at) and then distribute the exponent. Seriously, I tried it and know for a fact my answer shouldn't be 1 x 10^31...

OpenStudy (vuriffy):

I don't believe your equation is correct. I am sure it is: d= Vi + 1/2at^2

OpenStudy (vuriffy):

Oh, I see your question, one second.

OpenStudy (wholock):

Yeah, that's the equation my teacher told me to use.

OpenStudy (vuriffy):

d = 0 + 1/2(20)(1)^2 = 10 d = 10 + 1/2(20)(2)^2 = 50 d = 50 + 1/2(20)(3)^2 = 140 d = 140 + 1/2(20)(4)^2 = 300 d = 300 + 1/2(20)(5)^2 = 550 Not sure if that's right, or if it should be 1 second every time.

OpenStudy (vuriffy):

d = 0 + 1/2(20)(1)^2 = 10 d = 10 + 1/2(20)(1)^2 = 20 d = 20 + 1/2(20)(1)^2 = 30 d = 30 + 1/2(20)(1)^2 = 40 d = 40 + 1/2(20)(1)^2 = 50

OpenStudy (wholock):

OMG! Thank you very much! I can take it from here. Thank you!

OpenStudy (vuriffy):

I hope it helps as much as possible. Medal me, if you wish! Anytime.

OpenStudy (wholock):

Thanks!